# Generalization Bounds via Meta-Learned Model Representations: PAC-Bayes and Sample Compression Hypernetworks

### v2

@@ Appendix G Algorithm and Architectures
Figure 7: The PAC-Bayes hypernetwork.
Figure 8: The Sample Compression hypernetwork.
Figure 9: The PAC-Bayes Sample Compression hypernetwork.
@@ Appendix H Numerical Experiment and Implementation Details
Pixels swap and binary MNIST and CIFAR100 experiments. Let MLP1 be the architecture of the feedforward network in the sample compressor, the message compressor, the encoder, and the reconstructor; MLP2 be the architecture of the feedforward network in the DeepSet module; MLP3 the architecture of the downstream predictor. We had the following components and values in our hyperparameter grid:
@@
More results on binary MNIST and CIFAR100 experiments.
Table 4: Comparison of different meta-learning methods and generalization bounds on the MNIST (CIFAR100) binary task. The 95% confidence interval is reported for generalization bound and test error, computed over 34 (50) test tasks.
@@
Table 4 depicts the various bounds presented in section 2 with their corresponding algorithms.

### v3

@@ Appendix H Numerical Experiment and Implementation Details
The code for all experiments is available at https://github.com/GRAAL-Research/DeepRM.
@@
Pixels swap and binary MNIST and CIFAR100 experiments.  Let MLP1 be the architecture of the feedforward network in the sample compressor, the message compressor, the encoder, and the reconstructor; MLP2 be the architecture of the feedforward network in the DeepSet module; MLP3 the architecture of the downstream predictor. We used the following components and values in our hyperparameter grid:
@@
Table 5: Comparison of different meta-learning methods and generalization bounds on the MNIST and CIFAR100) binary task. The 95% confidence interval is reported for generalization bound and test error, computed over the test tasks.
@@
More results on binary MNIST and CIFAR100 experiments.  We present in Figure LABEL:fig:decomposition_pixel_swap and Table 6 (Figure LABEL:fig:decomposition_cifar_100 and Table 7) the contribution of each of the terms impacting the bound value for a few algorithms on the MNIST 200 pixels swap (binary CIFAR100) task. In the figures, the cumulative contributions are displayed, while in the tables, the marginal contributions are displayed. The bounds are decomposed as follows:
  • The observed meta train “error”;
  • The “confidence penalty”, which corresponds to the term $-\ln\theta$ in Theorem 2.1 and the corresponding term in other bounds;
  • The “complexity term”, which corresponds to the KL factor in the PAC-Bayes bounds. The latter is further decomposed into the compression set probability and the message probability in our sample compression-based bounds.
When considering the decomposition on the 200 pixels swap experiment, we see that our approaches, despite having a larger error term, relies on a small message probability and an empty compression set to yield competitive bounds. In contrast, for Pentina & Lampert (2014), the complexity term profoundly impacts the bound, making it non-competitive. As for the decomposition on the CIFAR100 experiments, it is interesting to see that the bound from Zakerinia et al. (2024) and the one from PB SCH have a similar decomposition, whereas SCH+, despite being penalized by the message probability, relies on a better treatment of its error and confidence term to obtain best bound of the four considered algorithms. This is empirical evidence of the tightness of our bounds compared to those of the runner-ups, all factors (error, confidence $\theta$, …) being kept equal, thanks to the non-linear comparator function $\Delta$ (see Theorem 2.1).
Label shuffle experiment.  As suggested by one ICML reviewer, we performed an additional experiment, based on the one in Amit & Meir (2018) (as well as Zakerinia et al. (2024)). The label shuffle experiment goes as follows: we start with the MNIST multiclass dataset, and we create each new task by performing random permutations of the label space. There are 30 training tasks of 1000 examples, and we evaluated the methods on 10 tasks of 100 samples. Table 8 presents the obtained results. We witness that, despite an extensive grid search over hyperparameters, the PB SCH approach fails to learn to generalize to the test tasks. We empirically observed severe overfitting on meta-train tasks. We hypothesize that this is due to the DeepSet architecture, which might not be powerful enough to encode the subtleties of the tasks at hand.
Table 6: Bound value decomposition for a few algorithms on the 200 pixels swap task.
Algorithm Error Confidence penalty Complexity Total
(Pentina & Lampert, 2014) 0.028 $\pm$ 0.002 0.085 $\pm$ 0.000 0.127 $\pm$ 0.028 0.240 $\pm$ 0.030
(Zakerinia et al., 2024) 0.042 $\pm$ 0.198 0.045 $\pm$ 0.000 0.021 $\pm$ 0.160 0.108 $\pm$ 0.037
Error Confidence penalty Message prob. Comp. set prob. Total
SCH+ 0.073 $\pm$ 0.017 0.015 $\pm$ 0.002 0.003 $\pm$ 0.001 $0$$\pm$ 0.000 0.091 $\pm$ 0.022
PB SCH 0.077 $\pm$ 0.016 0.030 $\pm$ 0.001 0.005 $\pm$ 0.005 $0$$\pm$ 0.000 0.112 $\pm$ 0.021
Table 7: Bound value decomposition for a few algorithms on the binary CIFAR100 task.
Algorithm Error Confidence penalty Complexity Total
(Pentina & Lampert, 2014) 0.485 $\pm$ 0.068 0.270 $\pm$ 0.000 0.046 $\pm$ 0.069 0.801 $\pm$ 0.001
(Zakerinia et al., 2024) 0.283 $\pm$ 0.120 0.133 $\pm$ 0.000 0.537 $\pm$ 0.195 0.953 $\pm$ 0.315
Error Confidence penalty Message prob. Comp. set prob. Total
SCH+ 0.264 $\pm$ 0.133 0.078 $\pm$ 0.010 0.374 $\pm$ 0.017 0.027 $\pm$ 0.003 0.745 $\pm$ 0.101
PB SCH 0.298 $\pm$ 0.098 0.155 $\pm$ 0.008 0.521 $\pm$ 0.045 $0$$\pm$ 0.000 0.974 $\pm$ 0.022
Table 8: Comparison of different meta-learning methods on the MNIST label shuffle binary task. The 95% confidence interval is reported for generalization bound and test error, computed over 10 test tasks. The best (smallest) result in each column is bolded.
Algorithm MNIST label shuffle
Bound ($\downarrow$) Test error ($\downarrow$)
(Pentina & Lampert, 2014) 2.376 $\pm$ 0.001 0.900 $\pm$ 0.589
(Amit & Meir, 2018) 0.542 $\pm$ 0.034 0.023 $\pm$ 0.062
(Guan & Lu, 2022) - kl 3.199 $\pm$ 0.372 0.023 $\pm$ 0.006
(Guan & Lu, 2022) - Catoni 0.536 $\pm$ 0.063 0.030 $\pm$ 0.001
(Rezazadeh, 2022) 12.95 $\pm$ 0.001 0.902 $\pm$ 0.660
(Zakerinia et al., 2024) 3.779 $\pm$ 0.079 0.165 $\pm$ 0.023
PB SCH 0.997 $\pm$ 0.005 0.792 $\pm$ 0.070
Bound on average over the decoder output.